EOC 16.29 Now assume that the quota is set at a level higher than the current level of imports. The resulting equilibrium price paid by consumers in the imported softwood lumber market will O ARise Fall O CNot change O D May rise or fallEOC 16.27 Once again assuming that the quota is set at a level lower than the current level of imports, the resulting equilibrium price received by the producer in the imported Canadian softwood lumber market wil Rise O B Fall C Not change O D Cannot be determinedEOC 16.22 If the price of tomatoes in the U.S. changes, what will happen? O AThe price will change by more than $.50 O B The price will change by less than $.50 The price will change by $.50 O D The price will stay the sameEOC16.18 Now suppose trade is opened between the two countries. What will be the equilibrium price in the world market for computers? Price per in 1000s $1 $2 $3 $4 $5 $6 Quantity Computer Demanded Supplied Quantity U.S 30 35 40 45 50 Ouatity Quantity Demanded Supplied U.S. Japan 50 40 30 20 10 Japan 90 80 70 60 50 40 50 60 65 70 75$1 0 в $2
Answer #1

1. Now assume the quota is set at a level higher than the current level of imports .The resulting equilibrium price paid by consumers in the imported softwood lumber market will fall.Because as the level of imports increases , now the consumers would pay less price .Hence, option(B) is correct.

2. Once again assumed that quota is set at a level lower than the current level of imports, the resulting equilibrium price received by the producer in the imported Canadian softwood lumber market will rise. Because as the level of imports fall, price rises. Hence,option(A) is correct.

3. The equilibrium price in the world market is at which import of U.S equals to export from Japan. At price of $3, U.S imports = (70-40)= 30 and Japan exports = (60-30)=30 . So, world equilibrium price=$3 . Hence,option(C) is correct.